Given the following information:

Revenues:                      $18 million                    Average Inventory:         $6 million          

Liabilities:                      $10 million                    Average A/R:                $2 million

Total Expenses:             $7 million                      Average Fixed Assets:     $10 million

Cost of Goods Sold:       $3 million                      Accounts Payables:         $4 million

Assume no other assets or liabilities exist beyond what is articulated above:

  1. Compute Net Profit Margin
  2. Compute Total Asset Turnover
  3. Compute Return on Equity (ROE)
  4. Compute Inventory Turnover
  5. How much equity would have to be swapped out for debt to increase ROE by 1% assuming that nothing else changes?
  6. What is the firm’s sustainable growth rate if dividends are equal to $0.5 million?
Answer #1

A Net profit margin = Net profit/revenues

Revenue = 18 million

COGS = 3 million

Total expenses = 7 million

Net profit = revenue - COGS - total expenses = 18 - 3 - 7 = 8 MILLION

In absence of any other expense, gross profit will be net profit.

Net profit margin = 8/18 = 0.4444 = 44.44%

B Total asset turnover = revenue / total assets

total assets = inventory + accounts rec'ble + fixed assets

= 6+2+10 = 18 million

total asset turnover = 18/18 = 1

for every dollar of company's asset, it generates $1 revenue.

C Return on Equity = Net income/shareholder's equity

Shareholder's equity = total assets - total liabilities

liabilities = 10 million , share holder's equity = 18 - 10 = 8 million

ROE = 8/8 = 1

D Inventory turnover = COGS / average inventory

average inventory = 6 million , COGS = 3 million

Inventory turnover = 3/6 = 0.5

Hope this helps!!